If N is normal in G and a belongs to G is of order O(a), prove that the order, m of Na in G/N is a divisor of O(a)


1
Expert's answer
2021-11-25T09:56:03-0500

We denote the identity element of "G"  by "e".

Since "N" is normal in "G, G\/N"  inherits a group structure from "G" whence

"Na^s=(Na)^s =N,"

So

"a^s\u2208N;"

We note we cannot have

"0\u2264n<s,"

for if (3) binds, since

"Na^n=(Na)^n=Ne=N\u27f9a^n\u2208N,"

the order of "Na" in "G\/N"  is then less than s, contrary to our hypothesis. So

"n\u2265s"

We now exploit the euclidean algorithm to write

"n=qs+r,"

where "0\u2264r<s;" then

"r=n\u2212qs"

and

"a^r=a^{n\u2212qs}=a^na^{\u2212qs}=ea^{\u2212qs}=a^{\u2212qs;}"

now since "a^s\u2208N ," and "N" is a subgroup of "G"

"a^{\u2212s}=(a^s)^{\u22121}\u2208N,"

whence

"a^{\u2212qs}=(a^{\u2212s})^q\u2208N;"

but by (8), (10) forces

"a^r\u2208N;"

but now "0<r<s" contadicts the fact that "s" is the order of "Na" in "G\/N" is "s" , since (11) yields

"(Na)^r=Na^r=N."

Thus "r=0" and hence "s\u2223n."